# Efficiently Building a Domain-Specific Large Language Model from Scratch: A Case Study of a Classical Chinese Large Language Model

### v1

@@
General-purpose large language models demonstrate notable capabilities in language comprehension and generation, achieving results that are comparable to, or even surpass, human performance in many language information processing tasks. Nevertheless, when general models are applied to some specific domains, e.g., Classical Chinese texts, their effectiveness is often unsatisfactory, and fine-tuning open-source foundational models similarly struggles to adequately incorporate domain-specific knowledge. To address this challenge, this study developed a large language model, AI Taiyan, specifically designed for understanding and generating Classical Chinese. Experiments show that with a reasonable model design, data processing, foundational training, and fine-tuning, satisfactory results can be achieved with only 1.8 billion parameters. In key tasks related to Classical Chinese information processing such as punctuation, identification of allusions, explanation of word meanings, and translation between ancient and modern Chinese, this model exhibits a clear advantage over both general-purpose large models and domain-specific traditional models, achieving levels close to or surpassing human baselines. This research provides a reference for the efficient construction of specialized domain-specific large language models. Furthermore, the paper discusses the application of this model in fields such as the collation of ancient texts, dictionary editing, and language research, combined with case studies.
@@ 1 Introduction
As a carrier of Chinese cultural heritage, Classical Chinese texts embody a wealth of linguistic and cultural knowledge. Given that literary expression in Classical Chinese has spanned thousands of years, its characters, vocabulary, grammar, and phonological systems have continuously evolved, with each era producing complex linguistic phenomena that pose significant challenges to comprehension of modern readers. In addition to linguistic difficulties, understanding Classical Chinese texts also relies on extra historical and cultural knowledge. For instance, ancient authors often used multiple variants to refer to people and events, favored implicit and nuanced expression, and frequently drew upon allusions to convey emotions or ideas. The dual challenges of language and culture not only complicate the learning and reading of Classical Chinese for contemporary readers but also constitute major obstacles in the collation of ancient texts and research in Classical Chinese information processing.
To better support the collation of ancient texts, the teaching of Classical Chinese, and digital humanities research, we have developed a large language model specifically designed for Classical Chinese information processing, named “AI Taiyan” (AI 太炎). This paper introduces the construction of this model, covering its architectural design, data processing, base training, and fine-tuning. We further explore its potential applications in areas such as ancient text collation, lexicography, and linguistic research through case studies.
@@ 2 Related Work
2.1 Classical Chinese Information Processing Technologies
In recent years, driven by the practical needs of ancient text collation and publication, Classical Chinese teaching, and digital humanities research, studies on Classical Chinese information processing technologies have become increasingly in-depth. These studies encompass a wide range of tasks, including punctuation, word segmentation, part-of-speech tagging, named entity recognition, relation extraction, word sense disambiguation, and translation between Classical and Modern Chinese.
@@ 2.1 Classical Chinese Information Processing Technologies
In summary, neural network models and pre-trained language model methods have been widely applied in the field of Classical Chinese information processing. However, these methods impose high requirements on the scale and quality of training data. In Classical Chinese processing scenarios, many tasks are characterized by being “low-resource” but “knowledge-intensive”, posing significant challenges for language resource development and model research. Existing studies have shown that mechanisms such as data augmentation and joint learning can effectively alleviate these issues  (?). Given the strong capabilities of large language models in multi-task and few-shot learning, building dedicated large language models for Classical Chinese holds great potential for enhancing the comprehensive capabilities of Classical Chinese information processing.
@@ 3.2 Pretraining
During pre-training, the model primarily acquires foundational linguistic knowledge through the task of next-word prediction. As a machine learning model, the parameter weights of a large language model are determined by fitting the training data; thus, the scale, quality, and diversity of the training data are critical determinants of the model’s language capabilities. Although large-scale models for Classical Chinese focus on the understanding and generation of ancient Chinese texts, their outputs are often intended for contemporary readers. For example, tasks such as Classical-to-Modern translation require the conversion of Classical Chinese texts into modern Chinese, so the base language model must be proficient in both ancient and modern Chinese expressions. Additionally, both simplified and traditional Chinese texts are needed for practical applications in Classical Chinese information processing.
@@ 3.3 Supervised finetuning
Segmentation and Punctuation 州城西南隅有黄鹤楼者 给上述文本添加标点。 (Add punctuation to the above text.) 州城西南隅,有黄鹤楼者。
Allusion Recognition 刘桢徒有气,管辂独无年。 识别文本中的典故。 (Identify the allusions in the text.) 刘桢徒有气,使用了典故 刘桢病,喻指卧病穷丑。 (“刘桢徒有气” uses the allusion “刘桢病” (Liu Zhen’s illness) refering to being bedridden and suffering in poverty and ugliness.)
Word Explanation 州城西南隅,有黄鹤楼者。 文本中的“隅”是什么意思? (What does “隅” mean in the text?) 角落。 (Corner.)
Translation 州城西南隅,有黄鹤楼者。 将上文翻译成白话文。 (Translate the above text into Modern Chinese.) 在鄂州城的西南角, 有一座楼叫黄鹤楼。 (At the southwest corner of Ezhou city, there is a tower called Yellow Crane Tower.)
@@ 5.2 Lexicography
Second, annotation errors arising from misinterpretations by editors are not uncommon in existing dictionaries, posing significant challenges for dictionary revision. The following two examples illustrate this issue  [^19] From 吴铭, 《汉语大词典》校札笔记: “進利”, “進資”..
@@ 6 Conclusion
This study addresses the characteristics of “low-resource” and “knowledge-rich” scenarios in Classical Chinese information processing tasks, and proposes a method for building a large language model for Classical Chinese from scratch. Firstly, based on the requirements of domain knowledge learning and the current state of available data, a compact large language model architecture is designed (52 layers, 1.8 billion parameters). Subsequently, through data processing, base model training, and finetuning, the “AI Taiyan” Classical Chinese large language model is developed. This model demonstrates strong capabilities in interpreting classical texts, supporting a range of challenging Classical Chinese understanding tasks such as sentence segmentation and punctuation, allusion recognition, word explanation, and translation between Classical and modern Chinese. It is compatible with both simplified and traditional Chinese scripts. Experimental results show that, compared with general large models and other domain-specific models, “AI Taiyan” exhibits significant advantages across multiple evaluation tasks, achieving performance close to or surpassing human baselines. In addition, this paper discusses the potential applications of it in assisting with classical text collation, dictionary compilation and revision, and linguistic research.

### v2

@@
General-purpose large language models demonstrate notable capabilities in language comprehension and generation, achieving results that are comparable to, or even surpass, human performance in many natural language processing tasks. Nevertheless, when general models are applied to some specific domains, e.g., Classical Chinese texts, their effectiveness is often unsatisfactory, and fine-tuning open-source foundational models similarly struggles to adequately incorporate domain-specific knowledge. To address this challenge, this study developed a large language model, AI Taiyan, specifically designed for understanding and generating Classical Chinese. Experiments show that with a reasonable model design, data processing, foundational training, and fine-tuning, satisfactory results can be achieved with only 1.8 billion parameters. In key tasks related to language processing of Classical Chinese such as punctuation, identification of allusions, explanation of word meanings, and translation between ancient and modern Chinese, this model exhibits a clear advantage over both general-purpose large models and domain-specific traditional models, achieving levels close to or surpassing human baselines. This research provides a reference for the efficient construction of specialized domain-specific large language models. Furthermore, the paper discusses the application of this model in fields such as the collation of ancient texts, dictionary editing, and language research, combined with case studies.
@@ 1 Introduction
As a carrier of Chinese cultural heritage, Classical Chinese texts embody a wealth of linguistic and cultural knowledge. Given that literary expression in Classical Chinese has spanned thousands of years, its characters, vocabulary, grammar, and phonological systems have continuously evolved, with each era producing complex linguistic phenomena that pose significant challenges to comprehension of modern readers. In addition to linguistic difficulties, understanding Classical Chinese texts also relies on extra historical and cultural knowledge. For instance, ancient authors often used multiple variants to refer to people and events, favored implicit and nuanced expression, and frequently drew upon allusions to convey emotions or ideas. The dual challenges of language and culture not only complicate the learning and reading of Classical Chinese for contemporary readers but also constitute major obstacles in the collation of ancient texts and research in natural language processing of Classical Chinese.
To better support the collation of ancient texts, the teaching of Classical Chinese, and digital humanities research, we have developed a large language model specifically designed for Classical Chinese, named “AI Taiyan” (AI 太炎)  [^1] https://t.shenshen.wiki. This paper introduces the construction of this model, covering its architectural design, data processing, base training, and fine-tuning. We further explore its potential applications in areas such as ancient text collation, lexicography, and linguistic research through case studies.
@@ 2 Related Work
2.1 Natural Language Processing in Classical Chinese
In recent years, driven by the practical needs of ancient text collation and publication, Classical Chinese teaching, and digital humanities research, studies on natural language processing of Classical Chinese have become increasingly in-depth. These studies encompass a wide range of tasks, including punctuation, word segmentation, part-of-speech tagging, named entity recognition, relation extraction, word sense disambiguation, and translation between Classical and Modern Chinese.
@@ 2.1 Natural Language Processing in Classical Chinese
In summary, neural network models and pre-trained language model methods have been widely applied in the field of Classical Chinese processing. However, these methods impose high requirements on the scale and quality of training data. In Classical Chinese processing scenarios, many tasks are characterized by being “low-resource” but “knowledge-intensive”, posing significant challenges for language resource development and model research. Existing studies have shown that mechanisms such as data augmentation and joint learning can effectively alleviate these issues  (?). Given the strong capabilities of large language models in multi-task and few-shot learning, building dedicated large language models for Classical Chinese holds great potential for enhancing the comprehensive capabilities of Classical Chinese.
@@ 3.2 Pretraining
During pre-training, the model primarily acquires foundational linguistic knowledge through the task of next-word prediction. As a machine learning model, the parameter weights of a large language model are determined by fitting the training data; thus, the scale, quality, and diversity of the training data are critical determinants of the model’s language capabilities. Although large-scale models for Classical Chinese focus on the understanding and generation of ancient Chinese texts, their outputs are often intended for contemporary readers. For example, tasks such as Classical-to-Modern translation require the conversion of Classical Chinese texts into modern Chinese, so the base language model must be proficient in both ancient and modern Chinese expressions. Additionally, both simplified and traditional Chinese texts are needed for practical applications in Classical Chinese processing.
@@ 3.3 Supervised finetuning
Segmentation and Punctuation 州城西南隅有黃鶴樓者 給上述文本添加標點。 (Add punctuation to the above text.) 州城西南隅,有黃鶴樓者。
Allusion Recognition 劉楨徒有氣,管輅獨無年。 識別文本中的典故。 (Identify the allusions in the text.) 劉楨徒有氣,使用了典故 劉楨病,喻指臥病窮醜。 (“劉楨徒有氣” uses the allusion “劉楨病” (Liu Zhen’s illness) refering to being bedridden and suffering in poverty and ugliness.)
Word Explanation 州城西南隅,有黃鶴樓者。 文本中的“隅”是什麼意思? (What does “隅” mean in the text?) 角落。 (Corner.)
Translation 州城西南隅,有黃鶴樓者。 將上文翻譯成白話文。 (Translate the above text into Modern Chinese.) 在鄂州城的西南角, 有一座樓叫黃鶴樓。 (At the southwest corner of Ezhou city, there is a tower called Yellow Crane Tower.)
@@ 5.2 Lexicography
Second, annotation errors arising from misinterpretations by editors are not uncommon in existing dictionaries, posing significant challenges for dictionary revision. The following two examples illustrate this issue  [^20] From 吳銘, 《漢語大詞典》校札筆記: “進利”, “進資”..
@@ 6 Conclusion
This study addresses the characteristics of “low-resource” and “knowledge-rich” scenarios in natural language processing tasks of Classical Chinese, and proposes a method for building a large language model for Classical Chinese from scratch. Firstly, based on the requirements of domain knowledge learning and the current state of available data, a compact large language model architecture is designed (52 layers, 1.8 billion parameters). Subsequently, through data processing, base model training, and finetuning, the “AI Taiyan” Classical Chinese large language model is developed. This model demonstrates strong capabilities in interpreting classical texts, supporting a range of challenging Classical Chinese understanding tasks such as sentence segmentation and punctuation, allusion recognition, word explanation, and translation between Classical and modern Chinese. It is compatible with both simplified and traditional Chinese scripts. Experimental results show that, compared with general large models and other domain-specific models, “AI Taiyan” exhibits significant advantages across multiple evaluation tasks, achieving performance close to or surpassing human baselines. In addition, this paper discusses the potential applications of it in assisting with classical text collation, dictionary compilation and revision, and linguistic research.